Let T : R4--> R4 be the linear transfoi-mation




defined by




T(x1, X2, X3, X4) = (— x2, X1, — x4, X3).




Check that T4 = I. Also, find the minimal




polynomial of T.

1
Expert's answer
2022-03-01T17:51:58-0500

T(x1, x2, x3, x4) =(−x2, x1, −x4, x3)

We should write in matrix firstly:

"T=
\begin{bmatrix}
 0 & -1 & 0 & 0 \\
 1 & 0 & 0 & 0 \\
 0 & 0 & 0 & -1 \\
 0 & 0 & 1 & 0 \\
\end{bmatrix} ;"


In order to show that T matrix does not have real eigen values, should calculate "(\lambda\times I-A)" matrix:

"(\lambda\times I-A)=
\lambda \times
\begin{bmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1 \\
\end{bmatrix} -" "\begin{bmatrix}
 0 & -1 & 0 & 0 \\
 1 & 0 & 0 & 0 \\
 0 & 0 & 0 & -1 \\
 0 & 0 & 1 & 0 \\
\end{bmatrix} ="


"\begin{bmatrix}
 \lambda & 0 & 0 & 0 \\
 0 & \lambda & 0 & 0 \\
 0 & 0 & \lambda & 0 \\
 0 & 0 & 0 & \lambda \\
\end{bmatrix} -
\begin{bmatrix}
 0 & -1 & 0 & 0 \\
 1 & 0 & 0 & 0 \\
 0 & 0 & 0 & -1 \\
 0 & 0 & 1 & 0 \\
\end{bmatrix} =" "\begin{bmatrix}
 \lambda & 1 & 0 & 0 \\
 -1 & \lambda & 0 & 0 \\
 0 & 0 & \lambda & 1 \\
 0 & 0 & -1 & \lambda \\
\end{bmatrix};"


And find det of "(\lambda\times I-A)" matrix:

"\lambda \times (-1)^{1+1}
\begin{bmatrix}
 \lambda & 0 & 0 \\
 0 & \lambda & 1 \\
 0 & -1 & \lambda \\
\end{bmatrix}+
1 \times (-1)^{1+2} \times" "\begin{bmatrix}
 -1 & 0 & 0 \\
 0 & \lambda & 1 \\
 0 & -1 & \lambda \\
\end{bmatrix}="


"\lambda \times (\lambda^3+ \lambda) -(-\lambda^2-1)=" "\lambda \times (\lambda^3+ \lambda) -(-\lambda^2-1)= \lambda^4+\lambda^2+\lambda^2+1=" "\lambda^4+2\lambda^2+1;"

And then will equate the polynomial formed by the eigen values to 0:

"\lambda^4+2\lambda^2+1=0"

"\lambda^2" is greater than or equal to 0. The left-hand side of the equation is greater than or equal to 1, but the right-hand side of the equation is 0. The equality cannot be true. So this equation doesn't have real solutions. Thus,the matrix T does not have real eigen values.